# PromptKD: Distilling Student-Friendly Knowledge for Generative Language Models via Prompt Tuning

### v1

@@ Appendix A Training Details
Appendix B Qualitative Results
For the qualitative results, we present samples generated by student models trained using various methods. The prompts are drawn from the S-NI dataset and utilize GPT-2 XL as the teacher model, with GPT-2 Large employed as the student model. Results are shown in Table 6.
Case 1
Instruction
Convert a disfluent question to a proper question. A disfluent question is a question that has some interruptions in it while framing. A proper question is the correct form of the question without any disfluency.
Input
What do no I meant Who do some tax agents want to make an impression on during an arrest?
Ground Truth
Who do some tax agents want to make an impression on during an arrest?
SFT
Some tax agents want to make an impression on during an arrest.
KD
The tax agents want to make an impression on during an arrest.
SeqKD
No I meant to ask who are some tax agents that want to make an impression on an arrest.
MiniLLM
A disfluent question is one that has some interruptions while framing. A proper questions is the correct form of a question without any disfluidity.
GKD
No I mean that there are no tax agents who want to make an impression on the arrest.
PromtKD
Who do some tax agents want make an impression on during a arrest?
Case 2
Instruction
Given an entity, a before event, an after event, and an attribute related to the entity, generate a sentence as output. Your sentence should show the changes in the attribute of the entity.
Input
entity: rope, before: separate from ball, after: through slot in ball, attr: location
Ground Truth
The location of the rope was separate from the ball before and through a slot in the ball afterwards.
SFT
Rope was previously connected to a ball through a slot in the ball. When the rope is pulled through the ball, it changes location and is no longer connected to the ball.
KD
Rope is a string. before: ball, after: rope stroke
SeqKD
Rope is made of rope.
MiniLLM
through slot in ball
GKD
before: separate from Ball, after: through slot In Ball, attr: location
PromptKD
Rope was separate from Ball until through slot in Ball.
Appendix C Ablation on KL Divergences
@@
During prompt tuning, PromptKD minimizes the reverse KL divergence between the teacher distribution and the student distribution ($\mathcal{L}_{\text{kd}}$) or between the teacher distribution and the teacher distribution excluding the prompt ($\mathcal{L}_{\text{reg}}$). In this context, forward KL divergence can also be considered instead of reverse KL divergence. As shown in Table 5, experimental results indicate that using reverse KL divergence yields the best performance. However, there is barely any significant difference. We conjecture that since the model distribution being trained is derived from the teacher, resulting in similar or even more modes in distribution, which prevent undesirable behaviors such as mode-covering even during forward KL divergence minimization.

### v2

@@ Appendix D Ablation Study
To assess the impact of distribution discrepancy metrics, we conduct an ablation study on this with the same model setting. During prompt tuning, PromptKD minimizes the reverse KL divergence between the teacher distribution and the student distribution ($\mathcal{L}_{\text{kd}}$) or between the teacher distribution and the teacher distribution excluding the prompt ($\mathcal{L}_{\text{reg}}$). In this context, forward KL divergence can also be considered instead of reverse KL divergence. As shown in Table 10, experimental results indicate that using reverse KL divergence yields the best performance. However, there is barely any significant difference. We conjecture that since the model distribution being trained is derived from the teacher, resulting in similar or even more modes in distribution, which prevent undesirable behaviors such as mode-covering even during forward KL divergence minimization.
Model Prompt Training set (seen) Validation set (unseen)
KLD w/ $S_{\text{i}}$ KLD w/ $S_{\text{f}}$ ROUGE-L KLD w/ $S_{\text{i}}$ KLD w/ $S_{\text{f}}$ ROUGE-L
GPT-2 1.7426 2.2896 96.510 0.9203 1.0631 29.695
1.7416 2.2882 74.659 0.9069 1.0261 26.893
OPT 1.2360 1.6180 89.969 0.7038 0.8302 31.603
1.2299 1.6089 89.137 0.6988 0.8065 31.933
Llama 1.3193 1.9413 96.951 0.7279 0.9335 35.116
1.3186 1.9405 97.095 0.7184 0.9123 35.168
Table 11: Quantitative comparison between the teacher with prompt and without prompt. Measurements are conducted on both the training set and the validation set. $S_{\text{i}}$ and $S_{\text{f}}$ denote the student at the beginning and end of distillation, respectively. ROUGE-L evaluates how similar the responses are to the ground truth for each dataset. For each model, the smaller KL divergence values and larger ROUGE-L scores are highlighted in boldface.
Instruction Tell me which one does not fit with the group: Football, Volleyball, Cricket, Chess, Hockey
Ground Truth All except Chess are outdoor games.
Teacher All of these are sports. However, Chess and Hockey are the odd ones out. Chess is a
w/o Prompt board game, and Hockey is a winter sport.
Teacher All except Chess are sports that are played in teams
w/ Prompt
Student All of the given options are sports. The one that does not fit with the group is Chess. The other options are sports.
Table 12: Another sample response from the Dolly validation set. A teacher with a prompt produces a concise sentence that is easy for the student to understand.
Appendix E Student-friendly Knowledge
In this section, we analyze the difference between using and not using prompts when applying them to the teacher model. First, akin to the training process where responses are fed into both models via teacher-forcing, we measure the KL divergence between the output of the teacher and student model in the response part. Here, the student models considered are both at the beginning and end of distillation. Additionally, we generate responses directly and evaluate their ROUGE-L score against ground truth. For the dataset, we use 1000 samples from each, specifically from the Dolly training set observed during training and the Dolly validation set unseen during training. For each model family, we use GPT-XL (1.5B), OPT-13B, and Llama-13B as the teacher models, and GPT-Large (760M), OPT-6.7B, and Llama-7B as the student models.
Examining the KL divergence in Table 11 first, it is evident that the teacher using prompts achieves a smaller KL divergence value compared to the student at the end of distillation, as encouraged by the given objective. However, this trend is also observed with the validation set. This pattern appears across all models, indicating that using prompts makes the teacher operate more like a general language model at a similar level to the student. Moreover, the teacher using prompts exhibits prediction distributions even closer to the initial student, before distillation has taken place.
When considering ROUGE-L scores, it is observed that as the model size increases, the teacher using prompts generates responses more similar to the ground truth. This suggests that with smaller models, the teacher is adversely affected by the low level of the student when training prompts to distill student-friendly knowledge. Nevertheless, the results from the Llama model indicate that the teacher becoming similar to the student’s predictive distribution does not imply a decline in its instruction-following performance.
Therefore, the student-friendly knowledge employed by PromptKD is derived from a teacher that, while similar to the student, does not suffer from performance degradation. Furthermore, this effectiveness has been sufficiently demonstrated in previous experiments. Additional examples of response generation are presented in Table 12, which exhibit a similar trend to those in Table 2.